CCD39 RON Test

OFF: Stages disabled, cube, lap, flowerpot

LittleJoe 39 temp 24.30 (internal 34)
Bin: 1 - Framerate  990 Hz (2500 kpix/sec) - RON (e-):   8.39 (estimated)
Bin: 2 - Framerate  990 Hz ( 800 kpix/sec) - RON (e-):   5.33 (estimated)
Bin: 2 - Framerate  625 Hz ( 380 kpix/sec) - RON (e-):   3.83 (estimated)
Bin: 2 - Framerate  400 Hz ( 380 kpix/sec) - RON (e-):   3.85 (estimated)
Bin: 3 - Framerate  200 Hz ( 380 kpix/sec) - RON (e-):   3.94 (estimated)
Bin: 4 - Framerate  300 Hz ( 335 kpix/sec) - RON (e-):   3.96 (estimated)
Bin: 4 - Framerate  200 Hz ( 335 kpix/sec) - RON (e-):   4.12 (estimated)
Bin: 4 - Framerate  105 Hz ( 335 kpix/sec) - RON (e-):   4.48 (estimated)

NTP servers on wfs/adsec

Updated ntp servers on both wfs and adsec to ntp1.lbt.as.arizona.edu and ntp restarted

On Sky Tests

Pisces15

First Preset, star too faint for TV


Second preset changed catalogue maq from 7 to 9.5, everything Fine

Autogain: gains: 2.58600 1.48700 0.965000

SR measurement, exp time 1acq Pisces_S, FeII
2011-06-20 09:37:45.170635: Pisces SR
PISCES exptime set to   1.0 seconds
Saving PISCES optloopdata...
Saved tracknum: 20110620_093745  (pisces=10 frames  ccd47=0 frames optical loop data=4000 frames)
Executed offset : x=0.000 y=-3.000 (mm)
Current position: x=0.000 y=-5.000 (arcsec)
Saved tracknum: 20110620_094040  (pisces=10 frames  ccd47=0 frames optical loop data=4000 frames)

Open loop
Saved tracknum: 20110620_095023  (pisces=10 frames  ccd47=0 frames optical loop data=4000 frames)

NOTE: Loop opened cleanly due to skip frames (bad seeing)

TrackNo RefStar Mag El Wind DIMM/OL Rec bin #mod freq gain mod nph AntiDrift SR filter exp #frames disturb SN notes
20110620_093346 PISCES15 9.6 74 - NaN 1.10 092604 1 400 990 2.6 1.5 1.0 3 88 ON 45.5 FeII 1.64 um 1000 1 ONSKY NO  
20110620_093745 PISCES15 9.6 73 - NaN 1.25 092604 1 400 990 2.6 1.5 1.0 3 89 ON 40.7 FeII 1.64 um 1000 10 ONSKY NO  
20110620_094040 PISCES15 9.5 72 - 1.19 1.09 092604 1 400 990 2.6 1.5 1.0 3 94 ON 0.0 FeII 1.64 um 1000 10 ONSKY NO Acq for DARK
20110620_095023 PISCES15 9.8 70 - 1.50 NaN 092604 1 400 990 -1.0 -1.0 -1.0 3 74 ON - FeII 1.64 um 1000 10 ONSKY NO OL acq.
In piu, sulla stellina compagna di queste imagine si fa 34.6% e 37.7% di SR.

olpsf_clpsf.png


Preset again same star, Acquire decided bin 2, preset:

Autogain: gains: 5.37600 1.42400 1.83464

Seeing 1.5 - 1.7
Saved tracknum: 20110620_100444  (pisces=4 frames  ccd47=0 frames optical loop data=4000 frames)
Executed offset : x=0.000 y=-3.000 (mm)
Current position: x=0.000 y=-5.000 (arcsec)
Saved tracknum: 20110620_100650  (pisces=4 frames  ccd47=0 frames optical loop data=4000 frames)

Seeing went up to 2.0!

TrackNo RefStar Mag El Wind DIMM/OL Rec bin #mod freq gain mod nph AntiDrift SR filter exp #frames disturb SN notes
20110620_100239 PISCES15 9.7 68 - 1.33 1.45 180534 2 145 991 5.4 1.4 0.9 3 348 ON 16.0 FeII 1.64 um 1000 1 ONSKY NO - SR may be in error
20110620_100444 PISCES15 9.8 67 - 1.58 1.87 180534 2 145 991 5.4 1.4 0.9 3 318 ON 12.1 FeII 1.64 um 3000 4 ONSKY NO - SR may be in error
20110620_100650 PISCES15 10.0 67 - 1.84 2.26 180534 2 145 991 5.4 1.4 0.9 3 253 ON 0.0 FeII 1.64 um 3000 4 ONSKY NO acq for DARK

----

We saw frame skipping due to bad seeing, there was also quite some drift, so we put back the aintidrift but with threshold = 4

Preset again TCS failed (guiding system)

Tried to close the loop manually, after a short time shell rip! (10:54:19) apparently right after an antidrift application (and stop fastlink). Housekeeper complained about the Vccp = 0.

TCS did not come back (failure with guiding camera).

-- JavierArgomedo - 20 Jun 2011
Topic revision: r8 - 11 Sep 2012, EnricoPinna
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback